4. A car travels up a hill at a constant speed of 40 km/h and returns down the hill at a constant speed of 60 km/h. (a) The displacement for the round trip is Hint: r ? (b) The average velocity for the round trip is Hint: -=- (c) The average speed for the round trip is Hint: average speedtotal distance km/h Ax kmh Assume the distance for one way is d. , -d 40, and 1,4 60 total time Joyce Jeong spring-2019 Modulc 2 Homework The coordinate of a particle in meters is given by x) 16 30r, where the time t is in (a) will the particle be momentarily at rest? When? (b) What is the speed of the particle at t 2.0s 5. 6. The position of an object moving along a straight line is given by s0)--30r+20r -10r. where x is in meters, r20, is in scconds. (a) The average acceleration between 0 and 4.0 s is ? Hint: At (b) The instantancous acceleration at - 10s is dv dt Why? (e ) Is this particle moving with a constant acceleration? Hint: Is aconstant? dh dr What is its

please help with
4(c)

5

6.
Answer #1

Answer 4) c)

Average speed= total distance/total time

Let distance of one way be d km.

T1=d/40 and t2= d/60

Total time=0.041 d

So, avg speed= 2d/0.041 d =47.9=50 km/hr (Ans)

=====

Acc to HOMEWORKLIB RULES, I can answer only one at a time. Good luck!:)